A 2 C charge travels with a velocity of ~v = (1.0ˆi+2.0ˆj+2.0 ˆk)m/s through a uniform magnetic field of (5.0ˆi + 4.0ˆj + 1.0 ˆk)T. Calculate the magnetic force exerted on the electron, in vector form.


1
Expert's answer
2021-07-06T08:22:40-0400

Let’s write the magnetic force that acts on the particle travelling with a velocity in a magnetic field :


"F=q(v\times B)."

Let's first calculate the cross product "\vec{v}\times \vec{B}":


"\vec{v}\times \vec{B}=\begin{vmatrix}
 \hat{i} & \hat{j} & \hat{k} \\
 v_x & v_y & v_z \\
 B_x & B_y & B_z
\end{vmatrix},""\vec{v}\times \vec{B}=\hat{i}\begin{vmatrix}
 v_y & v_z \\
 B_y & B_z
\end{vmatrix}-\hat{j}\begin{vmatrix}
 v_x & v_z \\
 B_x & B_z
\end{vmatrix}+\hat{k}\begin{vmatrix}
 v_x & v_y \\
 B_x & B_y
\end{vmatrix},""\vec{v}\times \vec{B}=(v_yB_z-B_yv_z)\hat{i}-(v_xB_z-B_xv_z)\hat{j}+(v_xB_y-B_xv_y)\hat{k}."

Substituting components of "v" and "B", we get:


"\vec{v}\times \vec{B}=(2-8)\dfrac{T\cdot m}{s}\hat{i}-(1-10)\dfrac{T\cdot m}{s}\hat{j}+(4-10)\dfrac{T\cdot m}{s}\hat{k},""\vec{v}\times \vec{B}=(-6\ \dfrac{T\cdot m}{s})\hat{i}+(9\ \dfrac{T\cdot m}{s})\hat{j}-(6\ \dfrac{T\cdot m}{s})\hat{k}."

Finally, we can find the magnetic force:


"F=2\ C\cdot[(-6\ \dfrac{T\cdot m}{s})\hat{i}+(9\ \dfrac{T\cdot m}{s})\hat{j}-(6\ \dfrac{T\cdot m}{s})\hat{k}],""F=(-12\ N)\hat{i}+(18\ N)\hat{j}-(12\ N)\hat{k}."